2. According to the US Bureau of Labor Statistics, there is a 5.2% probability that a randomly selected employed individual h 2 and 3 please
Answer #1

ans 2)since this is the problem of conditional probability .

Let  the event that an individual has more than one job A

and the event that an individual is male B.

we have given P(A) = 0.052 and P(B | A) = 0.504.

now P(A and B) = P(B | A) * P(A) = 0.504 * 0.052 = 0.0262 = 2.262%.

ans 3) since this is the case of binomial distribution

P(effective) = 0.75 ; P(not effective) = 0.25
P(at least one not effective) = 1 - P(none are not effective)
= 1- P(all are effective) = 1 - (0.75)^12 = 0.968